ax/\n-1/(bx/\n+c)

Grade:

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Hello student,
Please confirm if the question is as follows:
\int \frac{\frac{ax}{n-1}}{\frac{bx}{n+c}}
If this is the question, the answer goes as follows:
\frac{n+c}{n-1}\int\frac{ax}{bx}dx =\frac{a}{b}\frac{n+c}{n-1} \int dx=\frac{a}{b}\frac{n+c}{n-1}x
Thanks & Regards
Sumit Majumdar,
askIITians Faculty
Ph.D,IIT Delhi